# On-chip microwave coherent source with in-situ control of the photon number distribution

### v2

@@ Abstract
Coherent photon sources are key elements in different applications, ranging from quantum sensing to quantum computing. In the context of circuit quantum electrodynamics, there have been multiple proposals for potential coherent sources of photons, but a well established candidate is still missing. The possibility of designing and engineering superconducting circuits behaving like artificial atoms supports the realization of quantum optics protocols, including microwave photons generation. Here we propose and theoretically investigate a new design that allows a tunable photon injection directly on-chip. The scheme is based on initiating a population inversion in a superconducting circuit that will act as the photon source of one or multiple target resonators. The key novelty of the proposed layout consists in replacing the usual capacitive link between the source and the target cavity with a tunable coupler, with the advantage of having on-demand control on the injected steady-state photons. We validate the dynamical control of the generated coherent states under the effect of an external flux threading the tunable coupler and discuss the possibility of employing this scheme also in the context of multiple bosonic reservoirs.
@@ I Introduction
Precise control of photon population in bosonic elements, such as superconducting cavities, holds significant importance from both fundamental and practical perspectives in the realm of quantum sensing and computing. [1, 2, 3, 4, 5, 6, 7, 8]. In this framework, a coherent photon source represents a key element to control the photon distribution in bosonic infrastructures [9, 10, 11, 12]. Photon generation within the optical spectrum can be realized through multiple methods involving atoms, quantum dots, and nonlinear crystals [13, 14, 15, 16]. In the context of Circuit Quantum ElectroDynamics (cQED), the majority of approaches take advantage of the atom-like nature of superconducting circuits and the strong dipole interaction between different elements [17, 18, 19]. The conventional approach to drive superconducting cavities is to have a capacitive connection to room temperature sources delivering microwave radiations, that generate specific coherent states depending on the quadrature components of the microwave pulse [9]
In this work, we present an alternative and versatile approach for generating a photon distribution with a distinct average number of steady-state photons directly in situ with an inherently coherent photon source. Our proposal is based on engineering the spectrum and the transition rates of an artificial Josephson atom, which is a practice that has been extensively studied in different cQED platforms [20, 21, 22, 23, 24]. In particular, we consider an artificial atom able to work under the masing regime. In this hypothesis, the artificial atom operates as a natural on-chip micromaser able to emit microwave photons coherently [25, 26, 27, 28, 29, 30]. When the artificial atom is coupled to a resonant cavity, serving as a reservoir, it generates a well defined number of steady-state photons within the cavity. The generated photon distribution depends on the transition rates between the energy levels protagonist of the masing dynamics, the loss rate in the target reservoir and the coupling between the photon source and the reservoir cavity [31]. Here we introduce a novel scheme for manipulating the coherent state generated within the target cavity. Instead of utilizing a fixed capacitive coupling between the photon source and the target cavity, we propose a tunable coupler scheme [32, 33]. This approach introduces tunability on the effective coupling strength that governs the interaction between the photon source and the target cavity, enabling precise control over the average number of injected photons. Moreover, assuming that the transition frequency of the photon source can also be tuned while maintaining the transition rates unaltered, we can extend the fan-out of the system to multiple reservoirs resonating at different frequencies. This extension allows to individually manipulate different reservoirs according to the same dynamics of the single reservoir case. Photons demultiplexing can be employed in many fields, from quantum computing [34, 35, 36] to continuous variable encoding [37, 38].
II Concept and implementation
A schematic of the setup for the implementation of the proposed system is shown in Fig. 1(a). The system consists of $\mathrm{N_{r}+2}$ elements, where $\mathrm{N_{r}}$ represents the number of bosonic reservoirs on the output. The latter have resonance frequencies equal to $\mathrm{\omega_{r_{i}}}$, where $\mathrm{i}$ ranges from 1 to $\mathrm{N_{r}}$ and $\mathrm{\omega_{r_{i}}\neq\omega_{r_{j}}}$ for any $\mathrm{i\neq j}$. The remaining two elements correspond to the tunable photon source and the tunable coupler, each modeled for simplicity as two-level systems with transition frequencies $\mathrm{\omega_{s}}$ and $\mathrm{\omega_{c}}$, that depend on the fluxes $\mathrm{\Phi_{s}}$ and $\mathrm{\Phi_{c}}$ respectively. By adequately engineering the system, the tunable coupler can be decoupled from the other elements, recreating an analog system of $\mathrm{N_{r}+1}$ elements. In this hypothesis, the effect of the coupler is to renormalize the characteristic frequencies and coupling rates of the system by an amount that depends on its frequency. The ability to flux-tune the resonance frequency of the coupler defines a dynamical knob that gives control over the interaction between the photon source and a specific target reservoir, as depicted in Fig. 1(b). In the following section, we review the physics underlying the masing regime, on which the scheme is built upon. Subsequently, we discuss the methodology to effectively engineer the proposed scheme in the case of a single target resonator. We outline the key ingredients to build an experimentally feasible system that enables precise control over the photon population in the resonator via an external flux controlling the frequency of the tunable coupler. In this framework, we discuss different strategies to prevent detrimental effects coming from flux noise and faulty fabrication processes. We asses all these properties through a study of the second-order correlation function denoted as $\mathrm{g^{(2)}(0)}$, which we use as an indicator of the emission performance of the photon source. We conclude the analysis of this scheme by simulating its behavior and discussing the extension to a scheme with multiple reservoirs, focusing on the capability to control each reservoir while avoiding cross-population individually.
@@ II Concept and implementation
Figure 1: (a) Sketch of the general setup with the corresponding spectral characteristics of each element. The first element from the left is the photon source. We ignore the higher energy levels required for population inversion and consider only the tunable transition energy $\mathrm{\hbar\omega_{s}(\Phi_{s})}$ that is set on resonance with the target reservoir. The middle element is the tunable coupler with energy $\mathrm{\hbar\omega_{c}(\Phi_{c})}$, capacitively linked to various reservoirs with capacity $\mathrm{C_{r_{i}c}}$ $\mathrm{i=1...N_{r}}$ and to the photon source with a capacitive coupling $C_{sc}$. (b) Depiction of the dynamics underlying photon injection. A counter-relaxation process represents population inversion in the system at a rate $\mathrm{\Gamma}$ from the ground state |$\mathrm{\ket{g}}$ to the excited state |$\mathrm{\ket{e}}$ of the system. The photon source is resonant with the cavity to populate it with several photons depending on the external flux $\mathrm{\Phi_{c}}$. (c) The average number of steady-state photons $\mathrm{n_{ph}}$ and the second-order correlation function $\mathrm{g^{(2)}(0)}$ in terms of the effective coupling strength $g_{\text{eff}}$ between the photon source and the cavity and in terms of the masing ratio $\Gamma\kappa/4g_{eff}^{2}$ where $\kappa$ is the loss rate of the reservoir.
II.1 Masing regime
Before introducing the central idea of this work, we revisit the main notions related to the dynamics of an artificial atom operating within the masing regime, upon which our photon control and generation scheme is built. Masing is a physical phenomenon that governs the dynamics of masers and occurs when stimulated emissions prevail over absorption processes. Traditional masers typically involve multiple atoms and operate within a regime characterized by weak coupling [39]. However, in the realm of cQED, the strong coupling achievable between an artificial Josephson atom and photons permits the observation of the masing regime even at the level of a single artificial atom. In recent years there have been multiple demonstrations of single atom masing in cQED, employing a variety of techniques [40, 41, 42, 43]. Some of these techniques are based on triggering population inversion directly on the single Josephson artificial atom [26, 20, 44, 45, 46], while others achieve it through hidden population inversion in dressed state systems [47, 28, 30]. All these approaches share a common requirement for observing the masing process: the transition frequency related to photon emission must be set on resonance with the frequency of the reservoir. To model the dynamics of a typical masing system, we adopt the treatment followed in Ref. [31]. We consider a reservoir characterized by a loss rate $\kappa$, which is capacitively coupled to an artificial atom. These two systems engage in resonant interaction via the transition frequencies characterized by population inversion in the artificial atom spectra. In the dynamics of the system, the condition of population inversion is represented by a counter relaxation rate $\Gamma$, which governs the driven transition from the low energy state |$\ket{g}$ to the high energy state |$\ket{e}$. By considering just the emitting transition to model the artificial atom coupled to the reservoir, the system dynamics is well described by the Jaynes-Cummings Hamiltonian [48]:
$\mathrm{\hat{H}_{\text{JC}}=\hbar\omega_{r}\hat{a}^{\dagger}\hat{a}+\frac{\hbar\omega_{s}}{2}\hat{\sigma}_{z}+\hbar g_{eff}(\hat{a}^{\dagger}\hat{\sigma}_{-}+\hat{a}\hat{\sigma}_{+})}\,.$ (1)
@@ II.1 Masing regime
where $\mathrm{\hat{a}^{\dagger}}$ and $\mathrm{\hat{a}}$ are the creation and annihilation operators for the cavity mode, $\mathrm{\hat{\sigma}_{z}}$ is the Pauli-Z operator, $\mathrm{\hat{\sigma}_{-}}$ and $\mathrm{\hat{\sigma}_{+}}$ are the lowering and raising operators for the artificial atom, $\mathrm{\omega_{r}}$ is the cavity frequency, $\mathrm{\omega_{s}}$ is the artificial atom transition frequency, and $\mathrm{g_{eff}}$ represents the coupling strength between the atom and the cavity. The resonant coupling establishes a photon-mediated interaction between the states |$\mathrm{\ket{N,e}\leftrightarrow\ket{N+1,g}}$. Here, |$\mathrm{\ket{N,g}}$ (|$\mathrm{\ket{N,e}}$) represents the system when the resonator is populated by $\mathrm{N}$ photons and the artificial atom is in the ground (excited) state. Assuming that the counter-relaxation process in the artificial atom transition is characterized by a rate $\mathrm{\Gamma\gg g_{eff}}$, the dominant transition becomes |$\mathrm{\ket{N,e}\rightarrow\ket{N+1,g}}$. This predominance is directly linked to the counter-relaxation rate overcoming the atom-cavity coupling strength, which makes the artificial atom consistently appear in the excited state from the reservoir perspective during the typical interaction time $\mathrm{(\sim 2\pi/g_{eff})}$. Consequently, we can assume that the artificial atom is initially in the excited state. In this hypothesis, if the cavity contains $\mathrm{N}$ photons, the atom-cavity coupling $\mathrm{g_{eff}\sqrt{N+1}}$ induces dynamic transitions between the states |$\mathrm{\ket{N,e}}$ and |$\mathrm{\ket{N+1,g}}$. Since $\mathrm{\Gamma\gg g_{eff}}$, the transition will take the form of an incoherent process where any population that starts to accumulate in the state |$\mathrm{\ket{N,g}}$ will quickly relax to the state |$\mathrm{\ket{N,e}}$. These two steps complete the transition from the state |$\mathrm{\ket{N,e}}$ to the state |$\mathrm{\ket{N+1,e}}$, which allows to populate the cavity. This process will inevitably stop when the coupling strength $g_{eff}\sqrt{N}$ becomes comparable or larger than the counter relaxation rate $\mathrm{\Gamma}$. In this last situation the transition between |$\mathrm{\ket{N,e}}$ and |$\mathrm{\ket{N+1,g}}$ corresponds to the typical coherent oscillations of dressed state systems. The scaling of the Fock state ladder can also be limited by other two phenomena: photon blockade [49, 50] and the quantum Zeno effect [51, 52]. The photon blockade effect primarily arises from the proportional relationship between the splitting of the transition energy levels $\mathrm{\Delta}$ and the number of photons in the cavity $\mathrm{(\Delta\propto\sqrt{N})}$. As the number of injected photons increases, the resonant transition progressively deviates from the resonance condition. It is worth noting that a technique proposed in [50], involving an auxiliary cavity linked to the artificial atom, can mitigate the photon blockade effect. Hence, we will not consider this effect in our subsequent analysis. The quantum Zeno effect, instead, occurs for large values of $\mathrm{\Gamma}$ and results from the decoherence associated with the pumping process, which affects the quantum behavior of the system and, in turn, the masing regime [52, 31]. Considering a system where these last two phenomena can be neglected, we study the condition in which the masing regime takes place. Following the model used above for describing the dynamics under the masing regime, we can define the loss rate from the cavity to be $\mathrm{\Gamma_{l}=N\kappa}$, while the emission rate linked to the atom-cavity transition will be $\mathrm{\Gamma_{e}=4Ng_{eff}^{2}/\Gamma}$ [31]. Connecting these rates, the emissions overcome the losses when the following condition is satisfied [31]:
$\frac{\Gamma_{l}}{\Gamma_{e}}=\mathrm{\frac{\Gamma\kappa}{4g_{eff}^{2}}<1}\,.$ (2)
@@
Under this condition the interaction between the artificial atom and the cavity leads to photon population of the cavity through the mechanism described before, allowing the observation of the masing regime. Moreover, using a mean-field approximation, it is possible to also extract an analytical expression of the average number of steady-state photons that can be stored in the cavity [31]:
$\mathrm{n_{ph}=\frac{\Gamma}{2\kappa}\left(1-\frac{\Gamma\kappa}{4g_{eff}^{2}}\right)}\,.$ (3)
@@
Eq. 3 shows that a system characterized by well-defined values of $\mathrm{\Gamma}$ and $\mathrm{\kappa}$ can attain a maximum number of photons, determined by the ratio $\mathrm{\Gamma/2\kappa}$, when the masing ratio is far from unity $\mathrm{(\Gamma\kappa/4g_{eff}^{2}\ll 1)}$.
II.2 Cavity control through tunable coupling

### v3

@@ Abstract
Coherent photon sources are key elements in different applications, ranging from quantum sensing to quantum computing. The possibility of designing and engineering superconducting circuits behaving like artificial atoms supports the realization of quantum optics protocols, including microwave photons generation. Here we propose and theoretically investigate a design that allows a tunable photon injection directly on-chip. Our approach enables control of the emission via an external knob while preserving, in principle, the stability and linewidth characteristics. This is achieved by replacing the conventional capacitive link between the source and the reservoir with a tunable coupler, with the advantage of avoiding direct dynamical manipulation of the photon source dynamics, providing pulsed control of the emission. We validate the dynamical control of the generated state under the effect of an external flux threading the tunable coupler and discuss the possibility of employing this scheme also in the context of multiple bosonic reservoirs.
@@ I Introduction
The generation of tailored electromagnetic radiation is the foundational operation behind the control of quantum elements, such as natural or artificial atoms [1, 2, 3, 4, 5]. Precise control over these signals is essential for the performance of quantum computation and sensing processes [6, 7, 8, 9, 10, 11, 12]. In superconducting architectures, electromagnetic radiation is typically generated by room-temperature microwave electronics and delivered to the chip via dedicated transmission lines [13, 14, 15]. However, this approach introduces several drawbacks. Connections between the quantum components and room-temperature electronics can introduce unwanted thermal and electromagnetic noise, reducing the coherence of quantum elements [16, 8]. Furthermore, such setups require numerous broadband connections with multiple stages of attenuation and filtering to mitigate noise at cryogenic stages [14]. As the number of connections scales linearly with the quantum elements, this results in a significant thermal load, necessitating larger refrigerators with increased cooling power [17]. These limitations indicate the need for alternative control strategies, particularly for large-scale quantum systems [18, 19, 20].
Cryogenic-integrated control electronics offer a promising alternative for addressing these challenges [21, 22, 23]. Performing most control operations directly at the cryogenic stage eliminates the need for direct connections to room-temperature electronics, thereby improving thermal isolation and enhancing the coherence of controlled quantum elements. In this context, on-chip single-atom lasers are particularly notable [24, 25, 26]. These cryogenic-integrated coherent sources exhibit unique features [27, 28] such as self-quenching and a pumping threshold for population inversion determined by circuit parameters rather than external pump power, unlike conventional multi-atom lasers [29, 30, 31, 32, 33]. Besides being interesting elements for the observation of different quantum optics phenomena, [34, 29, 27, 28, 30, 32, 35, 31], these sources also present exceptional performance in terms of phase noise showing linewidth close to the standard quantum limit [34, 36, 35, 37, 33, 21, 38]. These features make them particularly well-suited as reference signals for microwave operations at cryogenic temperatures [39] and for driving quantum systems such as qubits and oscillators in quantum computing applications [40, 21]. Furthermore, their narrow linewidth enables high high-frequency resolution, making these sources highly desirable for applications in spectroscopy, communication, and sensing [41, 42, 43, 44].
For practical implementation, precise control over the power and frequency of emitted radiation is crucial. In most experimental realizations, the photon emitter is embedded in a reservoir, and its output power is controlled by direct manipulation of the population inversion dynamics [32, 45, 33, 21]. The emission frequency is typically tuned by modulating the energy associated with the transition characterized by population inversion [33], while output power is regulated by varying the pumping rate that drives the population inversion dynamics [32, 46].
An illustrative example is single-atom maser based on Josephson junctions [32, 33, 21]. Here, power tunability is achieved via an external DC voltage that determines the rate of Cooper pair breaking, which is the mechanism responsible for population inversion [32]. However, emission control through this approach can impact the spectral features of the emissions [21]. Fast and dynamical control of the output power is required to implement these cryogenic sources for the control of quantum elements, like qubits or cavities [40]. In this scenario, the direct control on the source and the finite and sharp time shape of control pulses directly for manipulating the inversion population dynamics can induce a broadening of the emissions, resulting in a degradation of the spectral stability and the control fidelity [33].
In this work, we address these limitations by proposing a versatile scheme based on the introduction of a tunable coupler [47, 46, 48] between the photon source and its output reservoir. By incorporating a tunable coupler between the source and the reservoir, our scheme allows emission control without affecting their spectral content and the source stability. Within this scheme, the emissions are controlled by modulating the coupling strength, which enables power control via an external knob, without directly interfering with the photon source. Such decoupled control allows to maintain and preserve emission stability and linewidth of these coherent sources, which are critical figures of merit for the successful implementation of these devices for practical applications in different quantum platforms [4, 49, 50, 51, 18, 52, 53]. Furthermore, assuming the photon source itself is frequency-tunable [33], the fan-out can be extended by inserting different elements resonating at different frequencies at the output, creating a practical microwave demultiplexer. Such a design eliminates the need for direct connections between the quantum elements on-chip and room-temperature electronics, enhancing thermal isolation and scalability for next-generation quantum devices.
II Results and discussions
II.1 Concept and implementation
A schematic of the setup for the implementation of the proposed system is shown in Fig. 1(a). The system consists of $\mathrm{N_{r}+2}$ elements, where $\mathrm{N_{r}}$ corresponds to the number of elements on the output, here represented as harmonic reservoirs. The latter have resonance frequencies equal to $\mathrm{\omega_{r_{i}}}$, where $\mathrm{i}$ ranges from 1 to $\mathrm{N_{r}}$ and $\mathrm{\omega_{r_{i}}\neq\omega_{r_{j}}}$ for any $\mathrm{i\neq j}$. The remaining two elements correspond to the tunable photon source and the tunable coupler. These two elements are considered as two-level systems with frequencies $\mathrm{\omega_{s}}(\Phi_{s})$ and $\mathrm{\omega_{c}}(\Phi_{c})$, respectively. Differently from the tunable coupler, the photon source possesses an inner counter-relaxation rate $\Gamma$, indicated in Fig. 1(b), representative of the population inversion dynamics from which the spontaneous emission from the source originates. By adequately engineering the system, the tunable coupler can be dispersively coupled to the other elements, recreating an analog system of $\mathrm{N_{r}+1}$ elements, similar to its implementation in quantum computation for the realization of two-qubit gates [46, 54]. In this scenario, the effect of the coupler is to renormalize the characteristic frequencies and coupling rates of the system by an amount that depends on its frequency $\mathrm{\omega_{c}}(\Phi_{c})$. The ability to modulate the frequency of the coupler defines a dynamical knob that gives control over the interaction between the photon source and a specific target reservoir, as depicted in Fig. 1(b).
The presented scheme can be easily implemented with typical elements of cQED. The tunable coupler corresponds to a flux tunable superconducting qubit with a frequency adequately far from the other elements to fall under the dispersive regime. The photon source can be any generic system capable of initiating population inversion between two energy levels in its spectra, showing the universality of the proposed scheme to any photon source based on masing. In the following section, we review the physics underlying the masing regime and explore feasible approaches to achieve population inversion using elements of cQED. Subsequently, we discuss the methodology to effectively engineer the proposed scheme in the case of a single target resonator. We outline the key ingredients to build an experimentally
@@ II.1 Concept and implementation
Figure 1: Concept and implementation sketches. (a) Sketch of the general setup with the corresponding spectral characteristics of each element. The first element from the left is the photon source. We ignore the higher energy levels required for population inversion and consider only the tunable transition energy $\mathrm{\hbar\omega_{s}(\Phi_{s})}$ that is set on resonance with the target reservoir. The middle element is the tunable coupler with energy $\mathrm{\hbar\omega_{c}(\Phi_{c})}$, capacitively linked to various reservoirs with capacitors $\mathrm{C_{r_{i}c}}$ $\mathrm{i=1...N_{r}}$ and to the photon source with a capacitive coupling $C_{sc}$. (b) Description of the dynamics underlying photon injection. A counter-relaxation process represents population inversion in the system at a rate $\mathrm{\Gamma}$ from the ground state |$\mathrm{\ket{g}}$ to the excited state |$\mathrm{\ket{e}}$ of the system. The photon source is resonant with the reservoir to populate it with several photons depending on the external flux $\mathrm{\Phi_{c}}$. The tunable effective coupling between the source and the reservoir is labeled $g_{eff}$.
feasible system that enables precise control over the photon population in the resonator via an external flux controlling the frequency of the tunable coupler. In this framework, we discuss different strategies to prevent detrimental effects coming from flux noise and faulty fabrication processes. We asses all these properties through a study of the normalized second-order correlation function denoted as $\mathrm{g^{(2)}(0)}$, which we use as an indicator of the emission performance of the photon source. We conclude the analysis of this scheme by simulating its behavior and discussing the extension to a scheme with multiple harmonic reservoirs, focusing on the capability to control each reservoir while avoiding cross-population individually.
II.2 Masing regime
Masing is a physical phenomenon that governs the dynamics of masers and occurs when emissions prevail over absorption processes. Traditional masers typically involve multiple atoms and operate within a regime characterized by weak coupling [55]. In the realm of cQED, the strong coupling achievable between an artificial Josephson atom and photons [56] enables the observation of the masing regime even at the level of a single artificial atom. There have been multiple demonstrations of single atom masing in cQED, employing a variety of techniques [32, 45, 33, 57, 58, 21]. All these techniques leverage on the atom-like behavior of superconducting circuits and the possibility to engineer their spectra. The first experimental realization of masing in cQED involved a voltage-biased charge qubit [32]. In this setup, population inversion was achieved by connecting a drain electrode to one arm of the tunnel junction. A sufficiently high voltage bias breaks Cooper pairs, triggering sequential single-electron tunneling events. Under proper charge offset, these events will trigger a population inversion mechanism that can be represented as a counter-relaxation rate $\Gamma$ between the first two energy levels. The counter-relaxation rate depends on the circuit parameters of the qubit which can be engineered according to the required output power from the source. Other experimental validations are based on the same phenomena but employ different circuital designs [45, 33, 21]. A second approach consists in employing a transmon qubit coupled to an auxiliary low-Q cavity [59]. In this scheme, population inversion is initiated between the first two energy levels of the transmon through a two-photon excitation that drives the qubit to the second excited state. Coupling the low-Q cavity to the second transition |$\ket{f}\leftrightarrow\ket{e}$ increases its decay rate allowing it to significantly exceed the decay rate of the |$\ket{e}\rightarrow\ket{g}$ transition. This creates a population inversion between the first two energy levels with a counter-relaxation rate $\Gamma$ that depends on the quality factor of the cavity and the coupling strength with the transmon qubit. Other approaches are also based on the coupling between the qubit and resonator but rather achieve population inversion through the engineering of the transition rate between dressed states [60, 61, 62].
In all mentioned methods, the photon source can always be modeled as a two-level system with an effective counter-relaxation rate $\Gamma$, as depicted in Fig. 1(b). The dynamics of a single-atom maser can thus be characterized by studying the emission of a two-level system with a counter-relaxation rate $\Gamma$ capacitively coupled to a resonant harmonic reservoir with a loss rate $\kappa$. The full system dynamics is well described by the Jaynes-Cummings Hamiltonian [63]:
$\mathrm{\hat{H}_{\text{JC}}=\hbar\omega_{r}\hat{a}^{\dagger}\hat{a}+\frac{\hbar\omega_{s}}{2}\hat{\sigma}_{z}+\hbar g_{eff}(\hat{a}^{\dagger}\hat{\sigma}_{-}+\hat{a}\hat{\sigma}_{+})}\,,$ (1)
@@ II.2 Masing regime
where $\mathrm{\hat{a}^{\dagger}}$ and $\mathrm{\hat{a}}$ are the creation and annihilation operators for the reservoir mode, $\mathrm{\hat{\sigma}_{z}}$ is the Pauli-Z operator, $\mathrm{\hat{\sigma}_{-}}$ and $\mathrm{\hat{\sigma}_{+}}$ are the lowering and raising operators for the artificial atom, $\mathrm{\omega_{r}}$ is the reservoir frequency, $\mathrm{\omega_{s}}$ is the artificial atom transition frequency and $\mathrm{g_{eff}}$ represents the coupling strength between the atom and the reservoir. The resonant coupling establishes a photon-mediated interaction between the states |$\mathrm{\ket{N,e}\leftrightarrow\ket{N+1,g}}$. Here, |$\mathrm{\ket{N,g}}$ (|$\mathrm{\ket{N,e}}$) represents the system when the resonator is populated by $\mathrm{N}$ photons and the artificial atom is in the ground (excited) state. Assuming that the counter-relaxation process in the artificial atom transition is characterized by a rate $\mathrm{\Gamma\gg g_{eff}}$, the dominant transition becomes |$\mathrm{\ket{N,e}\rightarrow\ket{N+1,g}}$. This predominance is directly linked to the counter-relaxation rate overcoming the atom-reservoir coupling strength, thus, during the typical interaction time $\mathrm{(\sim 2\pi/g_{eff})}$, the reservoir interacts with the artificial atom in its excited state |$\mathrm{\ket{e}}$. Consequently, we can assume that the artificial atom is initially in the excited state. In this hypothesis, if the reservoir contains $\mathrm{N}$ photons, the coupling $\mathrm{g_{eff}\sqrt{N+1}}$ induces dynamic transitions between the states |$\mathrm{\ket{N,e}}$ and |$\mathrm{\ket{N+1,g}}$. Since $\mathrm{\Gamma\gg g_{eff}}$, the transition will take the form of an incoherent process where any population that starts to accumulate in the state |$\mathrm{\ket{N,g}}$ will quickly relax to the state |$\mathrm{\ket{N,e}}$. These two steps complete the transition from the state |$\mathrm{\ket{N,e}}$ to the state |$\mathrm{\ket{N+1,e}}$, which allows to populate the reservoir. This process will inevitably stop when the coupling strength $g_{eff}\sqrt{N}$ becomes comparable or larger than the counter relaxation rate $\mathrm{\Gamma}$. In this last situation the transition between |$\mathrm{\ket{N,e}}$ and |$\mathrm{\ket{N+1,g}}$ corresponds to the typical coherent oscillations of dressed state systems.
The scaling of the Fock state ladder can also be limited by two other phenomena: photon blockade [64, 59] and the quantum Zeno effect [65, 66]. The photon blockade effect primarily arises from the proportional relationship between the splitting of the transition energy levels $\mathrm{\Delta}$ and the number of photons in the reservoir $\mathrm{(\Delta\propto\sqrt{N})}$. As the number of injected photons increases, the resonant transition progressively deviates from the resonance condition. It is worth noting that a technique proposed in [59], involving an auxiliary reservoir linked to the artificial atom, can mitigate the photon blockade effect. Hence, we will not consider this effect in our subsequent analysis. The quantum Zeno effect, instead, occurs for large values of $\mathrm{\Gamma}$ and results from the decoherence associated with the pumping process, which affects the quantum behavior of the system and, in turn, the masing regime [66, 67]. Considering a system where these last two phenomena can be neglected, we study the condition for which the masing regime takes place. Following the model used above for describing the dynamics under the masing regime, we can define the loss rate from the reservoir to be $\mathrm{\Gamma_{l}=N\kappa}$, while the emission rate linked to the atom-reservoir transition will be $\mathrm{\Gamma_{e}=4Ng_{eff}^{2}/\Gamma}$ [67]. Connecting these rates, the emissions overcome the losses when the following condition is satisfied [67]:
$\lambda=\frac{\Gamma_{l}}{\Gamma_{e}}=\mathrm{\frac{\Gamma\kappa}{4g_{eff}^{2}}<1}\,,$ (2)
@@
where $\mathrm{\lambda}$ corresponds to the masing ratio, a fundamental parameter that regulates the emissions in a masing photon source. Under this condition the interaction between the artificial atom and the reservoir leads to photon population of the reservoir through the mechanism described before, allowing the observation of the masing regime. Moreover, using a mean-field approximation, it is possible also to extract an analytical expression of the average number of steady-state photons that can be stored in the reservoir [67]:
$\mathrm{n_{ph}=\frac{\Gamma}{2\kappa}\left(1-\lambda\right)}\,.$ (3)
@@
Eq. 3 shows that a system characterized by well-defined values of $\mathrm{\Gamma}$ and $\mathrm{\kappa}$ can attain a maximum number of photons, determined by the ratio $\mathrm{\Gamma/2\kappa}$, when the masing ratio is far from unity $\mathrm{(\lambda\ll 1)}$.
II.3 Control through tunable coupling
The model described above reveals that the reservoir population is strongly dependent on the specific choices of the parameters $\mathrm{\Gamma}$, $\mathrm{\kappa}$, and $\mathrm{g_{eff}}$ in a single-atom maser. Consequently, the tunability of one of these parameters provides complete control over the steady-state photon population in the reservoir.